# Binary Classification of Light and Dark Time Traces of a Transition Edge Sensor Using Convolutional Neural Networks

### v1

@@ III-A Training process
III-B Performance evaluation
The performance of the model is evaluated against the rest of the dataset that was not dedicated for training as illustrated in Fig. 3, consisting of 2898 light pulses and 7872 dark pulses. With our main interest towards the use of the CNN in the ALPS II experiment, we follow the approach of our previous work (Ref. [37]) and evaluate the performance of the model during hyperparameter optimization based on the detection significance given by [54, 55]
@@ III-B Performance evaluation
where $T_{\mathrm{obs}}=518\,\mathrm{h}$ is the observation time of the experiment (as used in Ref. [37]), $n_{\mathrm{s}}=2.8\cdot 10^{-5}\,\mathrm{Hz}$ is the assumed signal (1064 nm photon) rate and $\epsilon_{d}=0.5$ is the pessimistically evaluated detection efficiency taking into account all losses associated with the experimental setup [37]. The only analysis method dependent parameters are the closely related background rate ($n_{\mathrm{b}}$) and analysis efficiency ($\epsilon_{\mathrm{a}}$). The $\epsilon_{\mathrm{a}}$ is simply calculated as the percentage of correctly classified light pulses (true positive). The $n_{\mathrm{b}}$ on the other hand is calculated from the number of misclassified dark pulses ($N_{\mathrm{mdp}}$, false positives). Since the total of 8872 extrinsic background pulses were measured over a time period of 2 d, the used testing set containing the subset of 7872 dark pulses effectively corresponds to $(7872/8872)\cdot 2\,\mathrm{d}\approx 1.77\,\mathrm{d}$ time period. The effective background rate can thus be estimated from the number of misclassified dark pulses (false positives) as $N_{\mathrm{mdp}}/1.77\,\mathrm{d}$. It should be pointed out that the $S$ score is a threshold ($\mathrm{Th.}\in[0,1]$ corresponding to dark–light, respectively) dependent metric. Consequently, all the reported values of $S$ in this manuscript have been obtained after optimizing the threshold to maximize its value.
While the $S$ score will be used to determine the optimal combination of hyperparameters in the following section, we will later also evaluate the trained CNNs using the $F_{1}$ score ($F_{1}=2\times\mathrm{Precision}\times\mathrm{Recall}/(\mathrm{Precision}+\mathrm{Recall})$ which balances between precision (TP/(TP+FP)) and recall (TP/(TP+FN)) thus making it a commonly used evaluation metric for biased datasets (TP=True Pos., FP=False Pos., FN=False Neg.). The $F_{1}$ score describes the true classification performance of the CNN better than the $S$ score by directly measuring how well the CNN is able to correctly classify the light pulses while avoiding the misclassification of dark pulses. Thus, we will utilize the $F_{1}$ score in particular in Section VI, where we study the nature and origins of the background pulses which the CNNs are struggling to classify correctly in more detail. The $F_{1}$ score is also a threshold dependent metric and its reported values in later sections have been obtained after the threshold optimization. It should be pointed out that the threshold optimization has to be done separately for the $S$ and $F_{1}$ scores.
@@
Hyperparameter Optimization range Optimum
Nb. of conv. layers 3–10 6
Nb. of filters 20–150 45
Kernel size 3–20 12
Dropout rate 0–0.2 0.18
Nb. of dense layers 1–10 3
Max nb. of neurons 100–300 188
Learning rate $10^{-5}$$10^{-3}$ $5.2\cdot 10^{-4}$
Epochs 5–20 10
Batch size 32–128 99
$\langle S\rangle=1.26\pm 0.16$
@@
TABLE I: The considered search space for the hyperparameter optimization of the CNN using 2000 iterations of random search. The activation functions of the convolutional and dense layers were fixed to tanh and ReLu, respectively. The presented optimum corresponds to the maximum obtained average detection significance $\langle S\rangle$ (see details in Section III-B) for an ensemble of 5 CNNs trained and evaluated with differently (randomly) divided training, validation and testing sets, while the weight initialization for the CNN was fixed. The values of $S$ were calculated using the optimal threshold that maximizes its value. The number of epochs was later increased to 20 (see Section IV)
@@
Figure 5: A schematic illustration of the CNN ensembles used in this study. (a) The ensemble used in Section IV where the CNNs are trained and evaluated with exactly the same datasets but the initialization of their weights differ between each other. (b) The ensemble used in Section V and VI. The model weights were fixed by seeding them equivalently to that of the CNN in the ensemble #1 that achieved the highest $S$ score (Eq. (3)). Only the dataset division into training and testing sets was randomized for this ensemble. (c) A single CNN trained with data divided equivalently to that of the CNN in the ensemble #2 whose percentage of misclassified dark pulses was closest to the average of the ensemble. This model was used in the latter part of Section VI.
III-C Hyperparameter Optimization
@@
The hyperparameters of the CNN architecture introduced in Section III are optimized by 2000 iterations of random search, i.e. by training a total of 2000 models using randomized combinations of hyperparameters and choosing the one with highest evaluation metrics. The search space for the considered hyperparameters is presented in Table I. In order to reduce the susceptibility of the evaluated performance towards the random division of the dataset into training and testing sets (Fig. 3), each iterated combination of hyperparameters is evaluated using 5 CNNs trained and tested with different, randomly divided, datasets as described in Sections III-A and III-B. The initial weights of the CNNs were fixed between the iterations of the random search.
@@
The evaluated average $S$ scores ($\langle S\rangle$) and their associated standard deviations ($\sigma_{S}$) as a function of trainable parameters in the CNN are illustrated in Fig. 4. The highest $S$ scores are clearly associated with CNNs with smaller number of trainable parameters, making the training process more efficient. We determine the optimal combination of hyperparameters for the CNN that maximizes $\langle S\rangle-\sigma_{S}$, under the constraint of limiting the maximum number of trainable parameters to $0.5\cdot 10^{6}$. The chosen optimal model has a total of 297,057 parameters and reached $\langle S\rangle=1.26\pm 0.16$. The associated hyperparameters are presented in Table I.
IV Fine-tuning optimized CNN
@@
Figure 6: Different learning curves associated with the CNN ensemble #1; (a) The average loss (binary cross-entropy) as a function of epochs for the whole ensemble of 10 CNNs. The dashed vertical line indicates the epochs=10 obtained from the initial hyperparameter optimization (see Table I). This was then increased to 20 in order to identify underfitting or overfitting from the associated figure. (b) The $F_{1}$ score as a function of training set size for the CNN associated with highest $F_{1}$ (also highest $S$) within the ensemble. The error bars correspond to the standard deviation associated with 3 statistical repetitions in the CNN training, between which the used training data was shuffled. The used testing data was kept constant throughout the process.
@@
Next, we will proceed investigating and fine-tuning the previously optimized CNN (Table I). In order to identify possible underfitting or overfitting, we increase the number of epochs from the optimized value of 10 up to 20. We train an ensemble of 10 CNNs with fixed datasets but randomly initialized weights. This CNN ensemble will be referred to as ensemble #1 (see Fig. 5(a)). We noticed that the performance of CNNs is susceptible towards the random initialization of its weights. This is manifested as a high standard deviation in the $\langle S\rangle=0.98\pm 0.22$ (average threshold: $\langle\mathrm{Th.}\rangle=0.98\pm 0.012$) for ensemble #1 (see Fig. 5(a)). While the weight initialization is internally stochastic, we have initialized the ensemble #1 weights in a deterministic manner via specified random seeds. This enables us to optimize the weight initialization of the CNN associated with the chosen seed for reproducible results. The observed sensitivity towards weight initialization is particularly common for CNNs associated with dropout layers. Similar observations have been made in other CNN based machine learning implementations [46]. The high values of optimal thresholds with respect to the $S$ score reflect the importance of decreasing the background rate at the cost of analysis efficiency (see Eq. (3)).
@@
We also evaluate ensemble #1 using the $F_{1}$ score (see III-B) reflecting the sole classification performance of the ensemble. We obtain $\langle F_{1}\rangle=0.96\pm 0.006$ ($\langle\mathrm{Th.}\rangle=0.63\pm 0.34$), indicating reduced conservativity in the classification of light pulses when compared with the $S$ score. This demonstrates the importance of threshold optimization for different metrics. While the high threshold values resulting in low background rates at the expense of analysis efficiency is preferable for the $S$ score, the best performance for the CNN, in terms of correctly classifying the light and the dark pulses, is obtained at much lower thresholds where the $F_{1}$ score is maximized.
@@
TABLE II: A summary of the evaluated performance of the CNN ensemble #2 (Fig. 5(b)) in terms of detection significance and $F_{1}$ score (Section III-B). The detection significance has been calculated using Eq. (3) with fixed observation time $T=518\,\mathrm{h}$, detection efficiency $\epsilon_{\mathrm{d}}=0.5$ and signal rate $\epsilon_{\mathrm{s}}=2.8\cdot 10^{-5}\,\mathrm{Hz}$ based on the current realistic limitations for the ALPS II experiment [37]. The reported background rate ($n_{\mathrm{b}}$) is calculated from the average percentage of mislcassified dark pulses (false positives) by dividing its value by the effective observation time for the extrinsic background as explained in Section III-B. The reported standard deviation for the avg. miscalssified dark pulses is slightly lower than the average value, but is rounded up.
@@
V Performance of the CNN
After the optimization of the CNN’s hyperparameters (Section III-C) together with additional optimization of number of epochs and weight initialization (Section IV), we now proceed in studying how well the CNN actually performs for its designated purpose as a binary classifier for the light and dark pulses. Here, it is important to note that the $S$- and $F_{1}$ scores evaluated for the trained CNN can be expected to be susceptible to the random division of the dataset into training, validation and testing sets. This is particularly true in the case where a small subset of mislabeled pulses (e.g. target value for a dark pulse = 1) exists. Consequently, we again train an ensemble of 10 CNNs and study its average performance on classifying the pulses. We will keep referring to this as ensemble #2 that is schematically illustrated in Fig. 5(b). Unlike for the ensemble #1 studied in the previous section, the weights of all the models in the ensemble #2 are initialized similarly based on the results of the previous Section IV. However, all the CNNs in ensemble #2 are trained and evaluated with different randomly divided datasets according to Fig. 3.
@@ V Performance of the CNN
Figure 7: The detection significance as a function of the cut range associated with the $\mathrm{Ph}_{\mathrm{FFT}}$ calculated based on the fitting parameters obtained in frequency domain analysis. The cut range is defined as $\mu_{\mathrm{m}}-n_{1}\cdot\sigma,\mu_{\mathrm{m}}+n_{1}\cdot\sigma]$, where $n_{1},n_{2}\in 0,1/3,\dots,3$ and the $\mu_{\mathrm{m}}\approx-16.33\,\mathrm{mV}$ and $\sigma\approx 1.25\,\mathrm{mV}$ are obtained by fitting a skewed Gaussian to the distribution of the peak heights for the light pulses. All of the detection significances were determined as the average of 5 $S$ scores calculated for randomly chosen testing sets.
V-A Comparison to Cut-Based Analysis
Next, we want to compare the above presented pulse analysis using the CNN ensemble #2 with the traditional (non-ML) cut-based analysis. In the cut-based analysis the pulse is classified based on its associated fitting parameters $\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$ and $\chi^{2}_{\mathrm{FFT}}$ introduced in Section II. In order for a pulse to be classified as light, all of the pulse parameters must simultaneously lie within $\mu_{\mathrm{m}}-3\cdot\sigma,\mu_{\mathrm{m}}+3\cdot\sigma]$ for the parameter specific $\mu_{\mathrm{m}}$ and $\sigma$. The only exception is with the $\mathrm{Ph}_{\mathrm{FFT}}$, for which the cut range $\mu_{\mathrm{m}}-n_{1}\cdot\sigma,\mu_{\mathrm{m}}+n_{2}\cdot\sigma]$ will be rigorously optimized for $n_{1},n_{2}\in 0,1/3,\dots,3$. These cut ranges in terms of $\mu_{\mathrm{m}}$ and $\sigma$ are determined based on the fits of skewed Gaussians to the parameter distributions.
However, in order to make the comparison between the performance of the CNN ensemble #2 and cut-based analysis fair, we determine the cut ranges for the associated pulse parameters based on 1000 randomly selected light pulses. In analogy to machine learning, determining the cut region corresponds to the training of the model. The rest of the light data together with the whole extrinsic dataset is then used as testing data to evaluate the $S$ score. It should be noted that the triggers used for determining the cuts (training) and evaluating the $S$ score (testing) correspond exactly to the pulses used for training and evaluating the CNNs. Thus, the used light pulses have been filtered as described in Section II.

### v2

@@ 3.1 Training process
3.2 Performance evaluation
The performance of the model is evaluated against the rest of the dataset that was not dedicated for training as illustrated in Fig. 4, consisting of 2898 light pulses and 7872 dark pulses. With our main interest towards the use of the CNN in the ALPS II experiment, we follow the approach of our previous work (Ref. [Meyer2024first]) and evaluate the performance of the model during hyperparameter optimization based on the detection significance given by [Bityukov1998new, Bityukov2000observability]
@@ 3.2 Performance evaluation
where $T_{\mathrm{obs}}=518\,\mathrm{h}$ is the observation time of the experiment (as used in Ref. [Meyer2024first]), $n_{\mathrm{s}}=2.8\cdot 10^{-5}\,\mathrm{Hz}$ is the assumed signal (1064 nm photon) rate and $\epsilon_{d}=0.5$ is the pessimistically evaluated detection efficiency taking into account all losses associated with the experimental setup [Meyer2024first]. The only analysis method dependent parameters are the closely related background rate ($n_{\mathrm{b}}$) and analysis efficiency ($\epsilon_{\mathrm{a}}$). The $\epsilon_{\mathrm{a}}$ is simply calculated as the percentage of correctly classified light pulses (true positive). The $n_{\mathrm{b}}$ on the other hand is calculated from the number of misclassified dark pulses ($N_{\mathrm{mdp}}$, false positives). Since the total of 8872 extrinsic background pulses were measured over a time period of 2 d, the used testing set containing the subset of 7872 dark pulses effectively corresponds to $(7872/8872)\cdot 2\,\mathrm{d}\approx 1.77\,\mathrm{d}$ time period. The effective background rate can thus be estimated from the number of misclassified dark pulses (false positives) as $N_{\mathrm{mdp}}/1.77\,\mathrm{d}$. It should be pointed out that the $S$ score is a threshold ($\mathrm{Th.}\in[0,1]$ corresponding to dark–light, respectively) dependent metric. Consequently, all the reported values of $S$ in this manuscript have been obtained after optimizing the threshold to maximize its value.
While the $S$ score will be used to determine the optimal combination of hyperparameters in the following section, we will later also evaluate the trained CNNs using the $F_{1}$ score ($F_{1}=2\times\mathrm{Precision}\times\mathrm{Recall}/(\mathrm{Precision}+\mathrm{Recall})$ which balances between precision (TP/(TP+FP)) and recall (TP/(TP+FN)) thus making it a commonly used evaluation metric for biased datasets (TP=True Pos., FP=False Pos., FN=False Neg.). The $F_{1}$ score describes the true classification performance of the CNN better than the $S$ score by directly measuring how well the CNN is able to correctly classify the light pulses while avoiding the misclassification of dark pulses. Thus, we will utilize the $F_{1}$ score in particular in Section 6, where we study the nature and origins of the background pulses which the CNNs are struggling to classify correctly in more detail. The $F_{1}$ score is also a threshold dependent metric and its reported values in later sections have been obtained after the threshold optimization. It should be pointed out that the threshold optimization has to be done separately for the $S$ and $F_{1}$ scores.
@@
Hyperparameter Optimization range 1 Optimum 1 Optimization range 2 Optimum 2 Nb. of conv. layers 3–10 6 4–7 7 Nb. of filters 20–150 45 20–60 40 Kernel size 3–20 12 5–15 7 Dropout rate 0–0.2 0.18 0.05–0.2 (step: 0.01) 0.07 Nb. of dense layers 1–10 3 3 (fixed) 3 Max nb. of neurons 100–300 188 188 (fixed) 188 Learning rate $10^{-5}$$10^{-3}$ $5.2\cdot 10^{-4}$ $5.2\cdot 10^{-4}$ (fixed) $5.2\cdot 10^{-4}$ Epochs 5–20 10 20 (fixed) 20 Batch size 32–128 99 99 (fixed) 99 $\langle S\rangle=1.26\pm 0.16$ $\langle S\rangle=1.24\pm 0.05$
@@
Table 1: The initial search space for the hyperparameter optimization (Optimization range 1) of the CNN using 2000 iterations of random search. The activation functions of the convolutional and dense layers were fixed to tanh and ReLu, respectively. The presented optimum (Optimum 1) corresponds to the maximum obtained average detection significance $\langle S\rangle$ (see details in Section 3.2) for an ensemble of 5 CNNs trained and evaluated with differently (randomly) divided training, validation and testing sets, while the weight initialization for the CNN was fixed. The values of $S$ were calculated using the optimal threshold that maximizes its value. After finding the initial optimum, the search space was narrowed down (Optimization range 2) and the alternative optimum (Optimum 2) was found using 5000 iterations of random search, covering 20.8% of the total search space. Narrowing down the search space did not result in improved $\langle S\rangle$ and consequently the combination of hyperparameters used for the rest of the study was fixed to the initial Optimum 1. The number of epochs was later increased to 20 (see Section 4). Continuous range
@@
Figure 6: A schematic illustration of the CNN ensembles (unweighted averaging ensembles) used in this study. (a) The ensemble used in Section 4 where the CNNs are trained and evaluated with exactly the same datasets but the initialization of their weights differ between each other. (b) The ensemble used in Section 5 and 6. The model weights were fixed by seeding them equivalently to that of the CNN in the ensemble #1 that achieved the highest $S$ score (Eq. (3)). Only the dataset division into training and testing sets was randomized for this ensemble. (c) A single CNN trained with data divided equivalently to that of the CNN in the ensemble #2 whose percentage of misclassified dark pulses was closest to the average of the ensemble. This model was used in the latter part of Section 6.
3.3 Hyperparameter Optimization
@@
The hyperparameters of the CNN architecture introduced in Section 3 are optimized by 2000 iterations of random search, i.e. by training a total of 2000 models using randomized combinations of hyperparameters and choosing the one with highest evaluation metrics. The search space for the considered hyperparameters is presented in Table 1; Optimization range 1. In order to reduce the susceptibility of the evaluated performance towards the random division of the dataset into training and testing sets (Fig. 4), each iterated combination of hyperparameters is evaluated using 5 CNNs trained and tested with different, randomly divided, datasets as described in Sections 3.1 and 3.2. The initial weights of the CNNs were fixed between the iterations of the random search.
@@
The evaluated average $S$ scores ($\langle S\rangle$) and their associated standard deviations ($\sigma_{S}$) as a function of trainable parameters in the CNN are illustrated in Fig. 5. The highest $S$ scores are clearly associated with CNNs with smaller number of trainable parameters, making the training process more efficient. We determine the optimal combination of hyperparameters for the CNN that maximizes $\langle S\rangle-\sigma_{S}$, under the constraint of limiting the maximum number of trainable parameters to $0.5\cdot 10^{6}$. The chosen optimal model has a total of 297,057 parameters and reached $\langle S\rangle=1.26\pm 0.16$. The associated hyperparameters are presented in Table 1Optimum 1.
While the search space associated with the above described hyperparamter optimization is huge when compared with the 2000 iterations of random search, one might argue that the performance of the CNN would be still limited by the non-optimal combination of hyperparameters. To address this, we have further performed fine-optimization by significantly narrowing down the search space. Considering the convolutional layers the most critical ones when it comes to the performance of the model, we fixed all the parameters associated with the dense layers to the previously found optima, together with learning rate and batch size. We fixed the number of epochs to 20, as we later found that it had room for improvement from the previously optimized value of 10 (see Section 4). Meanwhile, we narrowed down the ranges of parameters associated with the convolutional layers so that they still include the previously found optima (see Table 1Optimization range 2). The size of the new search space is 24,000 and we optimized the associated hyperparameters via 5000 iterations of random search, covering a total of 20.8% of the whole search space. This coverage is generally considered adequate for finding a near-optimum combination of hyperparameter [Bergstra2012random]. Regardless, the found optimum is associated with $\langle S\rangle=1.24\pm 0.05$ aligning with the initial random search within the uncertainty. Thus, we conclude that the initially found optimal combination of hyperparameters must be close to optimum and the choice of hyperparameters is unlikely to significantly limit the performance of the CNN. Thus, we will proceed with the initially optimized model (Table 1; Optimum 1).
4 Fine-tuning optimized CNN
@@
Figure 7: Different learning curves associated with the CNN ensemble #1; (a) The average loss (binary cross-entropy) as a function of epochs for the whole ensemble of 10 CNNs. The dashed vertical line indicates the epochs=10 obtained from the initial hyperparameter optimization (see Table 1). This was then increased to 20 in order to identify underfitting or overfitting from the associated figure. (b) The $F_{1}$ score as a function of training set size for the CNN associated with highest $F_{1}$ (also highest $S$) within the ensemble. The error bars correspond to the standard deviation associated with 3 statistical repetitions in the CNN training, between which the used training data was shuffled. The used testing data was kept constant throughout the process.
@@
Next, we will proceed investigating and fine-tuning the previously optimized CNN (Table 1; Optimum 1). In order to identify possible underfitting or overfitting, we increase the number of epochs from the optimized value of 10 up to 20. We train an ensemble of 10 CNNs (unweighted averaging ensemble) with fixed datasets but randomly initialized weights. This CNN ensemble will be referred to as ensemble #1 (see Fig. 6(a)). We noticed that the performance of CNNs is susceptible towards the random initialization of its weights. This is manifested as a high standard deviation in the $\langle S\rangle=0.98\pm 0.22$ (average threshold: $\langle\mathrm{Th.}\rangle=0.98\pm 0.012$) for ensemble #1 (see Fig. 6(a)). While the weight initialization is internally stochastic, we have initialized the ensemble #1 weights in a deterministic manner via specified random seeds. This enables us to optimize the weight initialization of the CNN associated with the chosen seed for reproducible results. The observed sensitivity towards weight initialization is particularly common for CNNs associated with dropout layers. Similar observations have been made in other CNN based machine learning implementations [Ismail2020inceptiontime]. The high values of optimal thresholds with respect to the $S$ score reflect the importance of decreasing the background rate at the cost of analysis efficiency (see Eq. (3)).
@@
We also evaluate ensemble #1 using the $F_{1}$ score (see 3.2) reflecting the sole classification performance of the ensemble. We obtain $\langle F_{1}\rangle=0.96\pm 0.006$ ($\langle\mathrm{Th.}\rangle=0.63\pm 0.34$), indicating reduced conservativity in the classification of light pulses when compared with the $S$ score. This demonstrates the importance of threshold optimization for different metrics. While the high threshold values resulting in low background rates at the expense of analysis efficiency is preferable for the $S$ score, the best performance for the CNN, in terms of correctly classifying the light and the dark pulses, is obtained at much lower thresholds where the $F_{1}$ score is maximized.
@@
Table 2: A summary of the evaluated performance of the CNN ensemble #2 (Fig. 6(b)) in terms of detection significance and $F_{1}$ score (Section 3.2). The detection significance has been calculated using Eq. (3) with fixed observation time $T=518\,\mathrm{h}$, detection efficiency $\epsilon_{\mathrm{d}}=0.5$ and signal rate $\epsilon_{\mathrm{s}}=2.8\cdot 10^{-5}\,\mathrm{Hz}$ based on the current realistic limitations for the ALPS II experiment [Meyer2024first]. The reported background rate ($n_{\mathrm{b}}$) is calculated from the average percentage of misclassified dark pulses (false positives) by dividing its value by the effective observation time for the extrinsic background as explained in Section 3.2. The reported confidence intervals correspond to standard deviations associated with the ensemble average. The reported standard deviation for the avg. misclassified dark pulses is slightly lower than the average value, but is rounded up.
@@
5 Performance of the CNN
After the optimization of the CNN’s hyperparameters (Section 3.3) together with additional optimization of number of epochs and weight initialization (Section 4), we now proceed in studying how well the CNN actually performs for its designated purpose as a binary classifier for the light and dark pulses. Here, it is important to note that the $S$- and $F_{1}$ scores evaluated for the trained CNN can be expected to be susceptible to the random division of the dataset into training, validation and testing sets. This is particularly true in the case where a small subset of mislabeled pulses (e.g. target value for a dark pulse = 1) exists. Consequently, we again train an ensemble of 10 CNNs (unweighted averaging ensemble) and study its average performance on classifying the pulses. We will keep referring to this as ensemble #2 that is schematically illustrated in Fig. 6(b). Unlike for the ensemble #1 studied in the previous section, the weights of all the models in the ensemble #2 are initialized similarly based on the results of the previous Section 4. However, all the CNNs in ensemble #2 are trained and evaluated with different randomly divided datasets according to Fig. 4.
@@ 5 Performance of the CNN
Figure 8: The detection significance as a function of the cut range associated with the $\mathrm{Ph}_{\mathrm{FFT}}$ calculated based on the fitting parameters obtained in frequency domain analysis. The cut range is defined as $\mu_{\mathrm{m}}-n_{1}\cdot\sigma,\,\mu_{\mathrm{m}}+n_{2}\cdot\sigma]$, where $n_{1},n_{2}\in 0,1/3,\dots,3$ and the $\mu_{\mathrm{m}}\approx-16.33\,\mathrm{mV}$ and $\sigma\approx 1.25\,\mathrm{mV}$ are obtained by fitting a skewed Gaussian to the distribution of the peak heights for the light pulses. All of the detection significances were determined as the average of 5 $S$ scores calculated for randomly chosen testing sets. The confidence interval of the $S$ score corresponds to standard deviation associated with the average.
5.1 Comparison to Cut-Based Analysis
Next, we want to compare the above presented pulse analysis using the CNN ensemble #2 with the traditional (non-ML) cut-based analysis. In the cut-based analysis the pulse is classified based on its associated fitting parameters $\tau_{\mathrm{rise}}$, $\tau_{\mathrm{decay}}$, $\chi_{\mathrm{Ph}}^{2}$, $\mathrm{Ph}_{\mathrm{FFT}}$ and $\chi^{2}_{\mathrm{FFT}}$ introduced in Section 2. In order for a pulse to be classified as light, all of the pulse parameters must simultaneously lie within $\mu_{\mathrm{m}}-3\cdot\sigma,\mu_{\mathrm{m}}+3\cdot\sigma]$ for the parameter specific $\mu_{\mathrm{m}}$ and $\sigma$. The only exception is with the $\mathrm{Ph}_{\mathrm{FFT}}$, for which the cut range $\mu_{\mathrm{m}}-n_{1}\cdot\sigma,\mu_{\mathrm{m}}+n_{2}\cdot\sigma]$ will be rigorously optimized for $n_{1},n_{2}\in 0,1/3,\dots,3$. These cut ranges in terms of $\mu_{\mathrm{m}}$ and $\sigma$ are determined based on the fits of skewed Gaussians to the parameter distributions.
However, in order to make the comparison between the performance of the CNN ensemble #2 and cut-based analysis fair, we determine the cut ranges for the associated pulse parameters based on 1000 randomly selected light pulses. In analogy to machine learning, determining the cut region corresponds to the training of the model. The rest of the light data together with the whole extrinsic dataset is then used as testing data to evaluate the $S$ score. It should be noted that the triggers used for determining the cuts (training) and evaluating the $S$ score (testing) correspond exactly to the pulses used for training and evaluating the CNNs. Thus, the used light pulses have been filtered as described in Section 2.